# A progress on the binary Goldbach conjecture

### v8

@@ 3.1. Application to the Binary Goldbach Conjecture
$=\frac{3n}{4}+O(\frac{n\log\log n}{\log n})$ (3.4)
as the main term. Now we analyze the error term in a similar manner. By virtue of the prime number theorem, we can write
$\pi(\frac{3n}{4})\log\log\pi(\frac{3n}{4})=(\frac{3n}{4\log n}+O(\frac{n}{\log^{2}n}))(\log\log(\frac{3n}{4\log n}+O(\frac{n}{\log^{2}n})))$ (3.5)
@@
$\log\log(\frac{3n}{4\log n}+O(\frac{n}{\log^{2}n}))$ $=\log(\log(\frac{3n}{4\log n})+\log(1+O(\frac{1}{\log n})))\ll\log\log n$ (3.6)
@@
$\ll\frac{n\log\log n}{\log n}$ (3.7)
and by combining (3.3) and (3.7), we obtain
@@
Now, by virtue of Lemma 3.7, we set $\Re(w_{1})$ to be a prime number and choose $\Re(w_{1})$ so that
$\frac{n}{4}<\Re(w_{1})$ $\leq(1+\frac{1}{\log^{3}\frac{n}{4}})(\frac{n}{4})$ (3.9)
for all $n\geq 358772$, then it implies that $\Re(z_{1})\lesssim\Re(w_{1})$. It is easy to see that
@@
for all $t\geq 4$ and $n>n_{o}$ for some fixed $n_{o}\in\mathbb{N}$. This completes the proof. ∎
We are now ready to prove the binary Goldbach conjecture for all even numbers greater than some $n_{0}\in\mathbb{N}$. This result provides an alternative solution to our first result and in very few instances adopts the proof technique in [11]. The benefit of the strong version of Bertrand’s postulate (Lemma 3.7) is good enough to verify the asymptotic version of the binary Goldbach conjecture using this version of the squeeze principle, which is a slight variation of the version that appears in the paper [11].
@@
There exist some $n_{o}\in\mathbb{N}$ such that every even number $n\geq n_{o}$ can be written as a sum of two prime numbers.
@@
We note that the above statement is equivalent to the statement that the cCoPs $\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{P}})$ are non–empty for all even $n\geq n_{o}$.
By remark 3.4 all cCoPs basing on $\mathbb{C}_{\mathbb{N}}$ with generators $\geq 2$ are non–empty. By virtue of Lemma 3.9 all cCoPs $\mathcal{C}^{o}(n)$ and $\mathcal{C}^{o}(n+4)$ with even generators $n\geq n_{o}$ fulfil the requirements of the special squeeze principle (Lemma 3.3) or the squeeze principle with $\mathbb{M}:=\mathbb{N}$ and $\mathbb{B}:=\mathbb{P}$. Hence for each such n there is always a non–empty cCoP $\mathcal{C}^{o}(n+2,\mathbb{C}_{\mathbb{P}})$. We start with $\mathcal{C}^{o}(n_{o})$ and $\mathcal{C}^{o}(n_{o}+4)$ and continue this procedure with $\mathcal{C}^{o}(n_{o}+k)$ and $\mathcal{C}^{o}(n_{o}+k+4)$ for all even $k\geq 2$. We verify that all cCoPs $\mathcal{C}^{o}(n_{o}+k+2)$ for even $k\geq 2$ ad infinitum are non–empty. ∎
@@ 4. The generalized squeeze principle and applications
Let $\mathbb{A}\subset\mathbb{N}$ then we call the set
@@
a multivariate circle of partition generated by $n\in\mathbb{N}$ with base regulators $\bigotimes_{i=1}^{h}\mathbb{A}_{i}$ the $h$-fold direct product of the sets $\mathbb{A}_{i}$. We call members of the multivariate circle of partitions multivariate points. We denote the weight of each points as $||[x_{i}]||:=x_{i}\in\mathbb{A}_{i}$ and for the corresponding weight set of the multivariate circle of partitions
$||\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A}_{i})||:=\{(x_{1},x_{2},\ldots x_{h})\in\bigotimes_{i=1}^{h}\mathbb{A}_{i}~{}|~{}\sum\limits_{i=1}^{h}x_{i}=n\}.$
We denote $\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}$ as an axis of the multivariate circle of partitions $\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A}_{i})$ if and only if $x_{i}\in\mathbb{A}_{i}$ for each $1\leq i\leq h$ and
@@
We say the axis points $[x_{i}]$ for each $1\leq i\leq h$ are axis residents. We do not view the axis as any different among other axis of the form $\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}$ up to the rearrangements of its residents points. In special cases where the points
$[x_{k}]\in\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A}_{i})$
is such that $hx_{k}=n$, then we call $[x_{i}]$ the center of the multivariate circle of partitions. If it exists, then we call it as a degenerated axis $\mathbb{L}_{[x_{k}]}$ in comparison to the real axes $\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}$, where not all of the weights $x_{i}$ can be equal. We denote the assignment of an axis $\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}$ to the multivariate CoP $\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A}_{i})$ as
$\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}~{}\hat{\in}~{}\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A}_{i})\mbox{ which means }[x_{1}],[x_{2}],\ldots,[x_{h}]\in\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A}_{i})$
@@
for a fixed $n\in\mathbb{N}$ with $x_{i}\in\mathbb{A}_{i}$ for each $1\leq i\leq h$ or vice versa and the number of real axes of the multivariate circle of partitions as
@@
In the special case where we fix $h=2$ and take $\mathbb{A}_{i}=\mathbb{A}\subset\mathbb{N}$, then we obtain the circle of partitions
@@
This structure was studied extensively in [10] in the case where one allows just two axis points on their axes. The squeeze principle is the statement
@@
Indeed it has found some unexpected applications to Goldbach-type problems and has been applied to study additive prime number problems requiring certain partitions into certain subsets of the positive integers. The power of this principle allows one to exhaust any interval of the form $[n,n+t]$ for $t\geq 4$ in way to conclude that all even numbers in this interval can be written as the sum of two prime numbers. Much more generally the set may be extended to a general subset of the integers and it mostly suffices to check if the underlying conditions of the principle are all satisfied in order to run this test. The squeeze principle has an alternate version which is also valid when one allows a complex base set. We restate it here without proof as a preliminary
@@
with $x_{i}\in\mathbb{A}$ for all $11\leq i\leq h-$ 1 and $x_{i}<x_{h}$ for all $11\leq i\leq h-$ 1 such that
$x_{h}>w:=\mathrm{\max}\{u\in||\mathcal{C}(n,\mathbb{N})||~{}|~{}u\in\mathbb{A}\}>n-\sum\limits_{i=1}^{h-1}x_{i}$
@@
for $x_{i}<x_{h}$ for all $11\leq i\leq h-$ 1. Clearly $w\in\mathbb{A}$ with $x_{i}\in\mathbb{A}$ for $11\leq i\leq h-$ 1 so that there exists an axis
@@
Let $\mathbb{A}\subset\mathbb{N}$ and suppose $\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A})\neq\emptyset$ for infinitely many $n\in\mathbb{H}\subseteq\mathbb{N}$. If for each $t\in\mathbb{N}$ such that $n,n+t$ are not consecutive in $\mathbb{H}$ there exists at least an axis
@@
with $x_{i}\in\mathbb{A}$ for all $11\leq i\leq h-$ 1 and $x_{i}<x_{h}$ for all $11\leq i\leq h-$ 1 such that
$x_{h}>w:=\mathrm{\max}\{u\in||\mathcal{C}(n,\mathbb{N})||~{}|~{}u\in\mathbb{A}\}>n-\sum\limits_{i=1}^{h-1}x_{i}$
@@
Suppose that $\mathbb{A}\subset\mathbb{N}$ and let $k$ be the smallest number in $\mathbb{H}\subseteq\mathbb{N}$ such that $\mathcal{C}(k,\bigotimes_{i=1}^{h}\mathbb{A})\neq\emptyset$, which means $k$ is the smallest number in $\mathbb{H}$ such that it can be written as the sum of $h$ elements in $\mathbb{A}\subset\mathbb{N}$. Let us choose $t_{o}\in\mathbb{N}$ such that $k,k+t_{o}$ are not consecutive in $\mathbb{H}$, then by the hypothesis there exists at least an axis
$\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}~{}\hat{\in}~{}\mathcal{C}(k+t_{o},\bigotimes_{i=1}^{h}\mathbb{N})$
with $x_{i}\in\mathbb{A}$ for all $11\leq i\leq h-$ 1 and $x_{i}<x_{h}$ for all $11\leq i\leq h-$ 1 such that
$x_{h}>w:=\mathrm{\max}\{u\in||\mathcal{C}(k,\mathbb{N})||~{}|~{}u\in\mathbb{A}\}>k-\sum\limits_{i=1}^{h-1}x_{i}.$
It follows from Lemma 4.3 there exists an $s\in\mathbb{N}$ with $k<s<k+t_{o}$ such that
@@
which means $s\in\mathbb{H}$ can be written as the sum of $h$ (not all) possibly distinct elements of $\mathbb{A}$. Let us consider the sub-intervals $[k,s]$ and $[s,k+t_{o}]$. If $k<s:=k+t_{1}$ are not consecutive in $\mathbb{H}$ then there exists at least an axis
@@
with $x_{i}\in\mathbb{A}$ for all $11\leq i\leq h-$ 1 and $x_{i}<x_{h}$ for all $11\leq i\leq h-$ 1 such that
$x_{h}>w:=\mathrm{\max}\{u\in||\mathcal{C}(k,\mathbb{N})||~{}|~{}u\in\mathbb{A}\}>k-\sum\limits_{i=1}^{h-1}x_{i}.$
It follows from Lemma 4.3 there exists some $u\in\mathbb{H}$ with $k<u<k+t_{1}$ such that
@@
which means $u\in\mathbb{H}$ can also be written as the sum of $h$ (not all) possibly distinct elements of $\mathbb{A}$. We can similarly iterate this argument on the intervals $[k,u],[u,s],[s,k+t_{o}]$ so far as there exists an element of $\mathbb{H}$ in any of the intervals. By virtue of the hypothesis the $t\in\mathbb{N}$ can be chosen arbitrarily such that $k,k+t$ are not consecutive in $\mathbb{H}$ and with the existence of at least an axis
$\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}~{}\hat{\in}~{}\mathcal{C}(k+t_{o},\bigotimes_{i=1}^{h}\mathbb{N})$
with $x_{i}\in\mathbb{A}$ for all $11\leq i\leq h-$ 1 and $x_{i}<x_{h}$ for all $11\leq i\leq h-$ 1 such that
$x_{h}>w:=\mathrm{\max}\{u\in||\mathcal{C}(k,\mathbb{N})||~{}|~{}u\in\mathbb{A}\}>k-\sum\limits_{i=1}^{h-1}x_{i}.$

### v9

@@ 4.1. Application to the Binary Goldbach Conjecture
(4.2) $=\frac{3n}{4}+O(\frac{n\log\log n}{\log n})$
as the main term. Now, we analyze the error term in a similar manner. Using the prime number theorem, we can write
(4.3) $\pi(\frac{3n}{4})\log\log\pi(\frac{3n}{4})=(\frac{3n}{4\log n}+O(\frac{n}{\log^{2}n}))(\log\log(\frac{3n}{4\log n}+O(\frac{n}{\log^{2}n}))).$
@@
(4.4) $\log\log(\frac{3n}{4\log n}+O(\frac{n}{\log^{2}n}))$ $=\log(\log(\frac{3n}{4\log n})+\log(1+O(\frac{1}{\log n})))\ll\log\log n$
@@
(4.5) $\ll\frac{n\log\log n}{\log n}.$
By combining (4.1) and (4.5), we obtain
@@
Now, by virtue of the lemma 4.7, we set $\Re(w_{1})$ to be a prime number and choose $\Re(w_{1})$ so that
(4.7) $\frac{n}{4}<\Re(w_{1})$ $\leq(1+\frac{1}{\log^{3}\frac{n}{4}})(\frac{n}{4})$
for all $n\geq 358772$. It implies that $\Re(z_{1})\lesssim\Re(w_{1})$. We deduce
@@
for all $t\geq 4$ and $n>n_{0}$ for some fixed $n_{0}\in\mathbb{N}$. This completes the proof. ∎
We are now ready to prove the binary Goldbach conjecture for all even numbers greater than some $n_{0}\in\mathbb{N}$. This result provides an alternative solution to our first result and in very few instances adopts the proof technique in [10]. The benefit of the strong version of the Bertrand postulate (Lemma 4.7) is good enough to verify the asymptotic version of the binary Goldbach conjecture using this version of the squeeze principle, which is a slight variation of the version that appears in the paper [10].
@@
There exist some $n_{0}\in\mathbb{N}$ such that every even number $n\geq n_{0}$ can be written as a sum of two prime numbers.
@@
We note that the above statement is equivalent to the statement that: the cCoPs $\mathcal{C}^{o}(n,\mathbb{C}_{\mathbb{P}})$ are non–empty for all even $n\geq n_{0}$.
By the remark 4.4 all cCoPs basing on $\mathbb{C}_{\mathbb{N}}$ with generators $\geq 2$ are non–empty. Using the lemma 4.9, all cCoPs $\mathcal{C}^{o}(n)$ and $\mathcal{C}^{o}(n+4)$ with even generators $n\geq n_{0}$ fulfill the requirements of the special squeeze principle (Lemma 4.3) or the squeeze principle with $\mathbb{M}:=\mathbb{N}$ and $\mathbb{B}:=\mathbb{P}$. Hence, for each such $n$ there is always a non–empty cCoP $\mathcal{C}^{o}(n+2,\mathbb{C}_{\mathbb{P}})$. We start with $\mathcal{C}^{o}(n_{0})$ and $\mathcal{C}^{o}(n_{0}+4)$ and continue this procedure with $\mathcal{C}^{o}(n_{0}+k)$ and $\mathcal{C}^{o}(n_{0}+k+4)$ for all even $k\geq 2$. We verify that all cCoPs $\mathcal{C}^{o}(n_{0}+k+2)$ for even $k\geq 2$ ad infinitum are non–empty. ∎
@@ 5. The generalized squeeze principle and applications
In this section, we generalize the squeeze principle introduced and developed in the previous section.
Definition 5.1.
Let $\mathbb{A}\subset\mathbb{N}$. We call the set
@@
a multivariate circle of partition generated by $n\in\mathbb{N}$ with base regulators $\bigotimes_{i=1}^{h}\mathbb{A}_{i}$ the $h$-fold direct product of the sets $\mathbb{A}_{i}$. We call members of the multivariate circle of partitions multivariate points. We denote the weight of each point by $x_{i}]||:=x_{i}\in\mathbb{A}_{i}$ and for the corresponding weight set of the multivariate circle of partitions
$\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A}_{i})||:=\{(x_{1},x_{2},\ldots x_{h})\in\bigotimes_{i=1}^{h}\mathbb{A}_{i}~|~\sum\limits_{i=1}^{h}x_{i}=n\}.$
We denote an axis of the multivariate circle of partition $\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A}_{i})$ by $\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}$ if and only if $x_{i}\in\mathbb{A}_{i}$ for each $1\leq i\leq h$ and
@@
We say that the axis points $x_{i}]$ for each $1\leq i\leq h$ are axis residents. We do not view the axes to be different among other axis of the form $\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}$ up to the rearrangements of its residents points. In special cases where the points
$x_{k}]\in\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A}_{i})$
is such that $hx_{k}=n$, we call $x_{i}]$ the center of the multivariate circle of partitions. If it exists, then we call it a degenerated axis $\mathbb{L}_{[x_{k}]}$ in comparison with the real axes $\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}$, where not all of the weights $x_{i}$ can be equal. We denote the assignment of an axis $\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}$ to the multivariate CoP $\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A}_{i})$ by
$\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}~\hat{\in}~\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A}_{i})$
which means
$x_{1}],[x_{2}],\ldots,[x_{h}]\in\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A}_{i})$
@@
for a fixed $n\in\mathbb{N}$ with $x_{i}\in\mathbb{A}_{i}$ for each $1\leq i\leq h$ or vice versa. We denote the number of real axes of the multivariate circle of partitions by
@@
In the special case where we fix $h=2$ and take $\mathbb{A}_{i}=\mathbb{A}\subset\mathbb{N}$, we obtain the circle of partitions
@@
This structure was studied extensively in [9] in the case where one allows only two axis points on their axes. The squeeze principle is the statement
@@
Indeed, it has found some unexpected applications to Goldbach-type problems and has been applied to study additive prime number problems requiring certain partitions into certain subsets of the positive integers. The power of this principle allows one to exhaust any interval of the form $[n,n+t]$ for $t\geq 4$ in way to conclude that all even numbers in this interval can be written as the sum of two prime numbers. In general, the set may be extended to a general subset of the integers and it mostly suffices to check if the underlying conditions of the principle are all satisfied in order to run this test. In the previous section, we proved:
@@
with $x_{i}\in\mathbb{A}$ for all $1\leq i\leq h-1$ and $x_{i}<x_{h}$ for all $1\leq i\leq h-1$ such that
[…]
@@
for $x_{i}<x_{h}$ for all $1\leq i\leq h-1$. Clearly $w\in\mathbb{A}$ with $x_{i}\in\mathbb{A}$ for $1\leq i\leq h-1$ so that there exists an axis
@@
Let $\mathbb{A}\subset\mathbb{N}$ and suppose that $\mathcal{C}(n,\bigotimes_{i=1}^{h}\mathbb{A})\neq\emptyset$ for infinitely many $n\in\mathbb{H}\subseteq\mathbb{N}$. If for each $t\in\mathbb{N}$ such that $n,n+t$ are not consecutive in $\mathbb{H}$ there exists at least an axis
@@
with $x_{i}\in\mathbb{A}$ for all $1\leq i\leq h-1$ and $x_{i}<x_{h}$ for all $1\leq i\leq h-1$ such that
[…]
@@
Suppose that $\mathbb{A}\subset\mathbb{N}$ and let $k$ be the smallest number in $\mathbb{H}\subseteq\mathbb{N}$ such that $\mathcal{C}(k,\bigotimes_{i=1}^{h}\mathbb{A})\neq\emptyset$, which means $k$ is the smallest number in $\mathbb{H}$ such that it can be written as the sum of $h$ elements in $\mathbb{A}\subset\mathbb{N}$. Let us choose $t_{0}\in\mathbb{N}$ such that $k,k+t_{0}$ are not consecutive in $\mathbb{H}$. By the hypothesis there exists at least an axis
$\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}~\hat{\in}~\mathcal{C}(k+t_{0},\bigotimes_{i=1}^{h}\mathbb{N})$
with $x_{i}\in\mathbb{A}$ for all $1\leq i\leq h-1$ and $x_{i}<x_{h}$ for all $1\leq i\leq h-1$ such that
[…]
It follows from the lemma 5.4 that there exists an $s\in\mathbb{N}$ with $k<s<k+t_{0}$ such that
@@
which means $s\in\mathbb{H}$ can be written as the sum of $h$ (not all) possibly distinct elements of $\mathbb{A}$. We consider the sub-intervals $[k,s]$ and $s,k+t_{0}]$. If $k<s:=k+t_{1}$ are not consecutive in $\mathbb{H}$, then there exists at least an axis
@@
with $x_{i}\in\mathbb{A}$ for all $1\leq i\leq h-1$ and $x_{i}<x_{h}$ for all $1\leq i\leq h-1$ such that
[…]
It follows from the lemma 5.4 that there exists some $u\in\mathbb{H}$ with $k<u<k+t_{1}$ such that
@@
which means $u\in\mathbb{H}$ can also be written as the sum of $h$ (not all) possibly distinct elements of $\mathbb{A}$. We can similarly iterate this argument on the intervals $k,u],[u,s],[s,k+t_{0}]$ in so far as there exists an element of $\mathbb{H}$ in any of the intervals. By virtue of the hypothesis, the $t\in\mathbb{N}$ can be arbitrarily chosen such that $k,k+t$ are not consecutive in $\mathbb{H}$ and with the existence of at least an axis
$\mathbb{L}_{[x_{1}],[x_{2}],\ldots,[x_{h}]}~\hat{\in}~\mathcal{C}(k+t_{0},\bigotimes_{i=1}^{h}\mathbb{N})$
with $x_{i}\in\mathbb{A}$ for all $1\leq i\leq h-1$ and $x_{i}<x_{h}$ for all $1\leq i\leq h-1$ such that
[…]